An Olympic class sprinter starts a race with an acceleration of 4.50 m/s2. What is her speed 2.40s later?


1
Expert's answer
2020-08-24T11:56:05-0400

"v=v_0+at=0+4.50\cdot2.40=10.8(m/s)"